Ana runs 4.00 m, 40.0° north of east;2.00 m, east; 2.00 m, east; 5.20 m, 30.0° south of west; and 6.50 m, south before stopping for water break. Find her resultant displacement from where she started.
1
Expert's answer
2020-09-16T10:10:02-0400

"x=4\cdot\cos40°+2+2-5.2\cdot\cos30°=2.56(m)"


"y=4\cdot\sin40°-5.2\cdot\sin30°-6.5=-6.53(m)"


"d=\sqrt{2.56^2+6.53^2}=7.01(m)"